OpenStudy (anonymous):

Find the sum of a finite geometric sequence from n = 1 to n = 6, using the expression −2(5)^n − 1

OpenStudy (anonymous):

\[{\rm S}=\frac{a_1(1-r^n)}{1-r}\]

OpenStudy (anonymous):

answer choices 1223 -1023 -7812 7812

OpenStudy (anonymous):

can you identify the common ratio ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, to find \(a_1\) plug in 1 for n

OpenStudy (anonymous):

1223

OpenStudy (anonymous):

what is the first term ?

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

-2(5)^(n-1) => -2(5)^(1-1) => -2(5)^0 => -2(1) => -2

OpenStudy (anonymous):

\[{\rm S_n}=\frac{a_1(1-r^{\ \rm n})}{1-r}=\frac{(-2)(1-5^6)}{1-5}\]

OpenStudy (anonymous):

-3125

OpenStudy (anonymous):

-7812